# Reflexive Guidance: Improving OoDD in Vision-Language Models via Self-Guided Image-Adaptive Concept Generation

### v1

@@ 5 Conclusion and Limitations
Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00352184).
@@ Appendix A Implementation Details
All experiments are implemented with Python 3.9 and PyTorch 1.9, using NVIDIA A100 80GB GPUs. For the InternVL2-llama3-76B model, due to its significant computational requirements, we employ 4 A100 GPUs with the VLLM (Kwon et al., 2023) library to reduce time overhead. For the remaining models, a single A100 GPU is sufficient to run the experiments.
@@ B.1 Dataset
We sample 25% from each dataset while maintaining the original proportion of samples for each label across the benchmarks.
@@ B.2 Comprehensive Experimental Results
Tab. B.2.1 and B.2.2 provide comprehensive results for experimental results on the ImageNet200 and CIFAR10 benchmark, respectively. For the CIFAR10 benchmark, we employ only GPT-4o as the proprietary model.
CLIP on GPT-4o valid query set. We evaluate OpenCLIP on inputs where GPT-4o generates valid responses to analyze whether GPT-4o’s superior results are influenced by input bias, such as the rejection of difficult inputs. OpenCLIP (GPT-4o valid) shows slightly better performance than the overall performance on the entire $25\%$ subset. However, despite this improvement, GPT-4o still outperforms OpenCLIP on both tasks. This demonstrates GPT-4o’s superior visual recognition capability and its ability to produce better confidence scores through its generated responses.
Scalability with image resolution. As shown in Tab. B.2.2, GPT-4o consistently outperforms OpenCLIP in both image recognition and OoDD tasks, particularly excelling in OoD distinguishability with an AUROC of $98.26\%$ and an FPR of $2.62\%$. However, GPT-4o returns the lowest number of valid responses among all compared models. QWEN processes low-resolution images better than high-resolution ones, although its overall performance remains poor. In terms of AUROC on OoD datasets, QWEN assigns higher confidence scores to ID classes for OoD inputs compared to ID inputs. Similarly, LLaVA-v1.6 processes low-resolution images better than high-resolution images. In contrast, GLM-4v performs worse on the CIFAR10 benchmark than on the ImageNet200 benchmark, including the valid response rate. One notable observation is that InternVL2-26B outperforms InternVL2-76B overall on the CIFAR10 benchmark.
Highly biased confidence scores. The compared models tend to produce OoD scores of either $0.0$ or $100.00$, as shown in Fig. 3. This indicates that the models typically classify input images as either ID or OoD with extremely high certainty, e.g., assigning $100.00$ to only one class and $0.00$ to all other classes. InternVL2-76B exhibits this tendency more strongly than GPT-4o. It produces almost no responses with OoD scores between $0.0$ and $100.00$ whereas GPT-4o generates some responses in this range.
@@
OpenCLIP 100.00 (23,031) 87.41 2.39 21.95 62.42 / 85.31 71.55 / 78.36 42.92 / 92.49 47.98 / 89.62 47.49 / 90.68 61.50 / 83.54
OpenCLIP (GPT-4o valid) 85.49 (19,690) 88.08 2.66 19.81 59.52 / 86.12 70.56 / 78.21 40.26 / 93.24 45.66 / 90.30 44.85 / 91.59 58.75 / 84.32
GPT-4o 85.49 (19,690) 89.78 2.42 18.64 22.30 / 92.08 38.94 / 81.41 2.06 / 97.58 7.45 / 95.85 3.78 / 97.17 23.50 / 88.50
@@
QWEN 72.81 (29,537) 15.90 76.96 705.77 100.00 / 51.91 81.80 / 28.91 99.63 / 28.02 78.39 / 53.73 96.77 / 27.95 96.98 / 28.21 92.01 / 31.01
InternVL2-4B 96.49 (39,143) 81.84 13.49 113.77 70.21 / 69.71 42.62 / 92.75 39.25 / 92.94 17.45 / 93.94 44.53 / 86.11 58.87 / 78.75 43.83 / 89.43
InternVL2-26B 99.80 (40,487) 96.22 1.87 24.47 22.65 / 87.72 0.27 / 98.91 0.00 / 99.04 1.35 / 98.37 12.01 / 93.03 10.95 / 93.61 4.77 / 96.66
InternVL2-76B 99.95 (40,549) 90.93 2.36 37.34 100.00 / 85.91 100.00 / 95.99 100.00 / 96.64 100.00 / 95.62 100.00 / 92.02 100.00 / 90.34 100.00 / 94.29
@@
In Sec. 3.6, we provide the reasoning results of GPT-4o and InternVL2-26B on the ImageNet200 benchmark. As shown in Tab. 2, GPT-4o demonstrates strong visual recognition abilities. It provides detailed descriptions of images, including aspects such as color, object shape, and relationships with other classes, leading to predictions in fine-grained categories. For example, GPT-4o correctly identifies the bald eagle image by describing its distinctive features, such as the white head, yellow beak, and large wingspan. For the image classified as lorikeet, GPT-4o provides highly detailed descriptions of the object’s visual features, including its vibrant green plumage, and arrives at its conclusion by comparing these features with other possible classes among the provided options, such as toucans or goldfinches. Due to its impressive visual recognition ability, GPT-4o’s misclassified predictions often involve classes closely related to the ID class, such as centipede and millipede.
@@
We additionally present the reasoning results of LLaVA-v1.6 on the same benchmark to analyze its low performance, despite reporting valid response rates comparable to other models. As shown in Tab. B.3.1, LLaVA-v1.6 shows a less detailed image interpretation compared to GPT-4o. It primarily predicts images into coarse-level categories, such as birds or flowers, and its explanations rely more on common sense than on visual features. This results in incorrect predictions, as evidenced in the cases where actual ID $\rightarrow$ prediction OoD. It suggests that better interpretability of visual features is crucial for enhancing OoDD capabilities.
Table B.3.1: Reasoning behind predictions of LLaVA-v1.6 on the ImageNet200 benchmark, excluding the case where the actual input is OoD but the prediction is ID. Due to the low valid response rate, none of LLaVA-v1.6’s responses corresponds to this case. Bold highlights the predicted class and its confidence score before applying softmax. Bold italic highlights notable points. Classes with a confidence score of $0.0$ are omitted.
@@ B.4 Implementation Details for Class Order
Table B.4.1: Grouping results on the ImageNet200 benchmark suggested by GPT-4o
@@ B.5 Detailed Results for Failure Cases
The number of each failure case. Tab. B.5.1 summarises the frequency of each failure case. Common failure cases can be categorized into six groups: 1) a mismatch between the predicted class and the class with the maximum confidence score, 2) failure to provide confidence scores for all classes, 3) responses that are not structured as the given format, 4) responses indicating an inability to identify the given image, 5) inclusion of classes not listed among the given, and 6) class duplication. For Case 2, we post-process the responses by assigning a confidence score of $0.0$ to the missing classes, as $200$ classes can be considered a substantial number to process effectively. Each failure case occurs uniformly across all datasets.
Table B.5.1: The number of each failure case on each model with the ImageNet200 benchmark
@@
Examples for inappropriate content. Tab. B.5.3 shows examples of images and their responses from each proprietary model. All inappropriate or harmful images we found are from SSB Hard. As discussed in Sec. 3.6, All three proprietary models decline to follow a given prompt if an input image contains such content, instead they return responses with “I’m sorry, I can’t help with that” or “I apologize, but I do not feel comfortable classifying or describing this image in detail”. Claude 3.5 Sonnet rarely generates responses corresponding to Case 4, but its frequency increases for the inappropriate classes in SSB Hard, such as undies or uzi. However, all open-source models generate predictions for them.
@@ B.6 Prompt Format
We provide the full prompt formats used for experiments in Sec. 3.5 and 3.6. In the prompt format, yellow, green, red, and blue boxes indicate a task description, an explanation of the reject class, guidelines, and examples for response format, respectively.
@@
Figure B.6.1: Prompt used for experiments in Sec. 3.5.
@@ B.6.2 Prompt for Reasoning
Figure B.6.2: Prompt used for reasoning in Sec. 3.6.
@@ B.6.3 Prompt for Class order
Figure B.6.3: Prompt used for GPT-4o to group classes for class order in Sec. 3.6.
@@ C.1 Suggested Class Post-processing
Despite the guidelines provided in Stage 1, responses that deviate from the guidelines are observed. These failure cases can be broadly categorized into four cases: 1) inclusion of ID class names, 2) failure to suggest exactly $N$ classes for each near- and far-OoD group, 3) duplicated classes, and 4) no suggestion due to failure in recognizing the image. We post-process the suggested answers to refine them into a feasible set of auxiliary OoD classes, except for Case 4. For Case 1, we exclude suggestions that are character-level identical to ID classes. For Case 2, if the total number of suggestions exceeds $2N$, we sample only $2N$ classes from the suggestions. If the total number of suggestions is less than $2N$, we use the suggestions as is without any supplementation. For Case 3, we remove one of the duplicated classes.
@@
InternVL2-26B 61.49 (2,564) 91.23 8.12 92.27 82.73 / 58.31 94.34 / 52.51 38.03 / 80.66 28.86 / 85.25 47.91 / 75.72 73.12 / 63.11
+ GPT-text 69.42 (2,895) 89.58 11.84 83.87 69.44 / 62.14 85.64 / 53.57 28.00 / 84.70 28.68 / 83.77 39.39 / 78.08 62.45 / 65.96
+ ReGuide 78.97 (3,293) 93.73 12.42 59.23 22.92 / 86.60 16.25 / 89.98 1.23 / 98.19 3.81 / 97.15 2.23 / 97.67 10.79 / 93.01
InternVL2-76B 99.63 (4,056) 89.09 5.93 48.20 100.00 / 77.83 100.00 / 70.43 100.00 / 96.65 100.00 / 91.80 100.00 / 93.35 100.00 / 80.70
+ReGuide 95.80 (3,995) 90.93 1.54 9.28 56.36 / 91.35 66.65 / 87.65 59.75 / 95.35 60.00 / 93.38 65.46 / 93.95 64.36 / 90.60
GPT-4o 87.58 (3,652) 90.08 3.30 7.94 11.90 / 94.41 33.69 / 85.05 2.03 / 97.95 6.03 / 94.48 3.63 / 97.51 18.76 / 91.08
+ ReGuide 79.42 (3,312) 91.50 1.55 8.66 25.12 / 96.73 37.66 / 92.52 23.47 / 97.05 32.60 / 95.93 28.26 / 96.81 32.57 / 94.57
@@
Tab. C.2.2 and C.2.3 present example responses of each models in Tab. 4. The first row in Tab. C.2.2 indicates the classes suggested by GPT-4o when ID class names are provided (i.e., GPT-text). The prompt used to obtain class suggestions of GPT-text can be found in Sec. C.5. A class name below each image is the ground truth of the corresponding image. The class suggestion set in GPT-text remains fixed regardless of the input images, as ID classes are static information, whereas ReGuide’s suggestions change based on the input image due to its image-adaptive nature.
The results on InternVL2-26B/-76B and GPT-4o demonstrate the effectiveness of ReGuide. Firstly, ReGuide suggests helpful classes for OoDD. For example, as shown in Tab. C.2.2, InternVL2+ReGuide suggests near-OoD classes for dotted image, such as polka dot, cherry red, and salmon-colored, which are highly close to its ground-truth. These suggested classes successfully guide the model to classify OoD inputs into one of the auxiliary OoD classes. For the softball image, the prediction of ReGuide is also closest to the ground-truth.
To ensure a fairer comparison between the models, we re-evaluate their performance based on the inputs that produce valid responses across all three models utilizing ReGuide. As shown in Table C.2.4, the effect of ReGuide is similar to that observed in Table C.2.1.
Table C.2.2: Example responses of InternVL2-26B on $5\%$ subset of ImageNet200 benchmark.
@@
Table C.2.3: Example responses of GPT-4o on $5\%$ subset of ImageNet200 benchmark.

### v2

@@ 5 Conclusion and Limitations
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) under Grant [RS-2024-00352184] and [RS-2024-00354675] funded by the Ministry of Science and ICT (MSIT).
@@ Appendix A Implementation Details
All experiments are implemented with Python 3.9 and PyTorch 1.9, using NVIDIA A100 80GB GPUs. For the InternVL2-llama3-76B model, due to the its significant computational requirements, we employ 4 A100 GPUs with the VLLM (Kwon et al., 2023) library to reduce time overhead. For the remaining models, a single A100 GPU is sufficient to run the experiments.
@@ B.1 Dataset
We sample 25% of each dataset, ensuring that the proportion of datasets in each benchmark are maintained. During sampling, we maintain the ratio of the number of samples for each label from the original dataset. Tables B.1.1 and B.1.2 present the number of images in each dataset for the ImageNet200 and CIFAR10 benchmarks, respectively. Fig. B.1.1 provides visual representations of this information. We additionally provide the class distributions of datasets for which label information is available. To improve clarity in visualization, we sampled 200 classes from SSB-Hard, which contains a large number of labels that would otherwise make the plot difficult to interpret. Fig. B.1.2 represents the class distributions for ImageNet200, NINCO, SSB-Hard, and Textures within the ImageNet200 benchmark, while Fig. B.1.3 represent the class distributions for CIFAR10, CIFAR100, MNIST, SVHN, Textures, and Places365 within the CIFAR10 benchmark. The lists of sampled images, along with the prompts and responses for each sample used in the main experiments are available at https://github.com/daintlab/ReGuide.
Table B.1.1: The number of images in each dataset in the ImageNet200 benchmark used in our experiments
Subset % ID Near-OoD Far-OoD Total
ImageNet200 NINCO SSB-Hard iNaturalist Textures Openimage-O
100% 9,000 5,879 49,000 10,000 15,869 5,160 94,908
25% 2,200 1,304 11,770 2,500 3,967 1,290 23,031
5% 400 249 1,970 500 793 258 4,170
Table B.1.2: The number of images in each dataset in the CIFAR10 benchmark used in our experiments
Subset % ID Near-OoD Far-OoD Total
CIFAR10 CIFAR100 MNIST SVHN Textures Places365 Tiny ImageNet
100% 9,000 9,000 70,000 26,032 5,640 35,195 7,793 162,660
25% 2,250 2,214 17,497 6,504 1,410 8,745 1,948 40,568
(a) 25% ImageNet200 benchmark
(b) 5% ImageNet200 benchmark
(c) 25% CIFAR10 benchmark
Figure B.1.1: The number of images in each dataset for each benchmark. Blue bars represent the full datasets, while red bars indicate the sampled datasets.
(a) ImageNet200
(b) NINCO
(c) SSB-Hard
(d) Textures
Figure B.1.2: The class distribution in each dataset for the ImageNet200 benchmark. Blue bars represent the full datasets, while red bars indicate the 25% sampled datasets.
(a) CIFAR10
(b) MNIST
(c) SVHN
(d) CIFAR100
(e) Places365
(f) Textures
Figure B.1.3: The class distribution in each dataset for the CIFAR10 benchmark. Blue bars represent the full datasets, while red bars indicate the 25% sampled datasets.
@@ B.2 Comprehensive Experimental Results
Tab. B.2.1 and B.2.3 provide comprehensive results for experimental results on the ImageNet200 and CIFAR10 benchmark, respectively. For the CIFAR10 benchmark, we employ only GPT-4o as the proprietary model.
CLIP on GPT-4o valid query set. We evaluate OpenCLIP on the inputs where GPT-4o generates valid responses to analyze whether GPT-4o’s better results are influenced by input bias, such as rejecting difficult inputs. As shown in Tab. B.2.1, OpenCLIP (GPT-4o valid) shows slightly better overall performance compared to its performance on the entire $25\%$ subset. Although there is an improvement, GPT-4o still outperforms OpenCLIP on both tasks. This demonstrates GPT-4o’s superior visual recognition capability and its ability to produce better confidence scores through its generated responses.
@@
OpenCLIP 100.00 (23,031) 87.41 2.39 21.95 62.27 / 85.31 71.48 / 78.36 42.76 / 92.49 47.83 / 89.62 47.47 / 90.68 61.42 / 83.54
OpenCLIP (GPT-4o valid) 85.49 (19,689) 88.08 2.66 19.81 59.33 / 86.12 70.31 / 78.20 40.10 / 93.24 45.66 / 90.30 44.60 / 91.59 58.54 / 84.32
GPT-4o 85.49 (19,689) 89.78 2.42 18.64 22.30 / 92.08 38.95 / 81.41 2.06 / 97.58 7.45 / 95.85 3.78 / 97.17 23.50 / 88.50
@@
Results on shared valid query set across LVLMs To enable a more rigorous comparison, we evaluate the models using the shared valid query set. Tab. B.2.2 presents the $25\%$ subset of the ImageNet200 results using this shared valid set. QWEN is excluded due to its exceptionally low valid ratio. The overall performance trends are consistent with those observed on the models’ individual valid query sets. GPT-4o outperforms all other compared models, and proprietary models demonstrate superior performance compared to open-source models, aligning with the results observed with their own valid query sets.
Table B.2.2: Comprehensive results on the shared valid query set of the ImageNet200 benchmark. For clarity, the full names of the compared models are provided in the footnotes below the table. $\downarrow$ and $\uparrow$ mean that lower and higher values are better, respectively. The AURC values are multiplied by $10^{3}$, and the other values are percentages. Bold represents the best performance among generative LVLMs.
Models ID Near-OoD Far-OoD All OoD
ImageNet200 NINCO SSB-Hard iNaturalist Textures Openimage-O
OpenCLIP 24.77 (5,707) 87.96 3.06 20.52 65.42 / 83.74 78.94 / 73.01 42.00 / 93.09 41.33 / 91.42 45.67 / 91.52 63.06 / 81.69
GPT-4o 91.74 2.79 14.86 18.31 / 91.49 44.96 / 77.05 1.62 / 97.97 6.09 / 95.72 2.56 / 97.54 25.69 / 86.44
Claude 3.5 Sonnet 87.84 7.96 75.22 58.64 / 68.24 83.51 / 55.44 7.20 / 95.31 9.04 / 94.81 18.93 / 89.34 51.95 / 71.99
Gemini Pro 1.5 90.48 2.93 23.71 25.76 / 86.51 62.37 / 69.57 0.88 / 98.18 6.27 / 95.10 4.88 / 96.69 35.45 / 82.09
LLaVA-v1.6 3.21 4.07 860.46 100.00 / 52.32 100.00 / 49.27 100.00 / 50.02 100.00 / 59.89 100.00 / 48.19 100.00 / 50.57
GLM-4v 69.04 2.94 81.79 100.00 / 77.70 100.00 / 68.14 100.00 / 82.98 100.00 / 82.65 100.00 / 82.80 100.00 / 74.92
InternVL2-26B 91.28 8.37 88.29 81.91 / 58.87 96.00 / 51.83 35.54 / 82.06 21.77 / 88.94 43.35 / 78.15 69.38 / 65.14
InternVL2-76B 90.02 6.42 83.00 61.69 / 67.38 83.07 / 56.69 2.20 / 97.12 9.78 / 93.33 13.92 / 91.26 50.44 / 73.00
    OpenCLIP-ViT-B-32, GPT-4o (2024-08-06), LLaVA-v1.6-Mistral-7B, GLM-4v-9B, InternVL2-Phi-3-mini-4B, InternVL2-InternLM2-Chat-26B, InternVL2-LLaMA3-76B
Scalability with image resolution. As shown in Tab. B.2.3, GPT-4o consistently outperforms OpenCLIP in both image recognition and OoDD tasks, particularly excelling in OoD distinguishability with AUROC of $98.26\%$ and FPR of $2.61\%$. However, GPT-4o returns the lowest number of valid responses among all compared models. QWEN processes low-resolution images better than high-resolution ones, though its overall performance remains poor. As shown in its AUROC on OoD datasets, QWEN assigns higher confidence scores to ID classes for OoD inputs than for ID inputs. LLaVA-v1.6 processes low resolution images better than high resolution images. However, GLM-4v performs worse on the CIFAR10 benchmark than on the ImageNet200 benchmark, including valid response rate. One notable observation is that InternVL2-26B outperforms InternVL2-76B overall on the CIFAR10 benchmark.
Highly biased confidence scores. The compared models tend to produce OoD score either $0.0$ or $100.00$, as shown in Fig. 4. This indicates that the models typically classify input images as either ID or OoD with extremely high certainty, e.g., assigning $100.00$ to only one class and $0.00$ to all other classes. InternVL2-76B exhibites this tendency more strongly than GPT-4o. It produces almost no responses with OoD scores between $0.0$ and $100.00$ whereas GPT-4o generates some responses in this range.
@@
QWEN 72.81 (29,537) 16.08 76.82 703.73 100.00 / 51.91 81.80 / 28.91 99.63 / 28.02 78.39 / 53.73 96.77 / 27.95 96.98 / 28.21 92.01 / 31.01
InternVL2-4B 96.49 (39,143) 81.84 13.49 113.77 70.21 / 69.71 42.62 / 92.75 39.25 / 92.94 17.45 / 93.94 44.53 / 86.11 100.00 / 78.75 58.87 / 89.43
InternVL2-26B 99.80 (40,487) 96.22 1.87 24.47 22.65 / 87.72 0.27 / 98.91 0.00 / 99.04 1.35 / 98.37 12.01 / 93.04 10.86 / 93.61 4.77 / 96.66
InternVL2-76B 99.95 (40,549) 90.93 2.36 37.34 100.00 / 85.91 100.00 / 95.99 100.00 / 96.64 100.00 / 95.65 100.00 / 92.02 100.00 / 90.34 100.00 / 94.29
@@
In Sec. 3.6, we provide the reasoning results of GPT-4o and InternVL2-26B on the ImageNet200 benchmark. As shown in Tab. 2, GPT-4o demonstrates strong visual recognition abilities. It provides detailed descriptions of images, including aspects such as color, object shape, and relationships with other classes, leading to predictions in fine-grained categories. For example, GPT-4o correctly identifies the bald eagle image by describing its distinctive features, such as the white head, yellow beak and large wingspan. For the image classified as lorikeet, GPT-4o provides highly detailed descriptions of the object’s visual features, including its vibrant green plumage, and arrived at its conclusion by comparing these features with other possible classes among the provided options, such as toucans or goldfinches. Due to its impressive visual recognition ability, GPT-4o’s misclassified predictions often involve classes closely related to the ID class, such as centipede and millipede.
@@
We additionally present the reasoning results of LLaVA-v1.6 on the same benchmark to analyze its low performance, despite reporting valid response rates comparable to other models. As shown in Tab. B.3.1, LLaVA-v1.6 shows much rough image interpretation than GPT-4o. It primarily predicts images into coarse-level categories, such as birds or flowers, and its explanations rely more on common sense than on visual features. This results in incorrect predictions, as evidenced in the cases where actual ID $\rightarrow$ prediction OoD. It suggests that better interpretability of visual features is crucial for enhancing OoDD capabilities.
Table B.3.1: Reasoning behind predictions of LLaVA-v1.6 on the ImageNet200 benchmark, excluding cases where the actual input is OoD but the prediction is ID. Due to the low valid response rate, none of LLaVA-v1.6’s responses corresponds to this case. Bold highlights the predicted class and its confidence score before applying softmax. Bold italic highlights notable points. We omit some classes, such as those with a confidence score of $0.0$.
@@ B.4 Implementation Details for Class Order
Table B.4.1: Groups and corresponding classes suggested by GPT-4o
@@ B.5 Detailed Results for Failure Cases
The number of each failure case. Tab. B.5.1 summarizes the frequency of each failure case. Common failure cases can be categorized into six groups: 1) a mismatch between the predicted class and the class with the maximum confidence score, 2) failure to provide confidence scores for all classes, 3) responses that are not structured as the given format, 4) responses indicating an inability to identify the given image, 5) inclusion of classes not listed among the given, and 6) class duplication. For Case 2, we post-process the responses by assigning a confidence score of $0.0$ to the missing classes, as $200$ classes can be considered a substantial number to process effectively. Each failure case occurs uniformly across all datasets.
Table B.5.1: The number of each failure case on each model with the 25% ImageNet200 benchmark
@@
Examples for inappropriate content. Tab. B.5.3 shows examples of images and their responses from each proprietary model. All inappropriate or harmful images we found are from SSB Hard. As discussed in Sec. 3.6, All three proprietary models decline to follow a given prompt if an input image contains such content, instead they return responses with “I’m sorry, I can’t help with that” or “I apologize, but I do not feel comfortable classifying or describing this image in detail”. Claude 3.5 Sonnet rarely generates responses corresponding Case 4, but its frequency increases for the inappropriate classes in SSB Hard, such as undies or uzi. However, all open-source models generate predictions for them.
@@
The effectiveness of guidelines and examples. We examine the effects of guidelines and examples to mitigate invalid responses using InternVL2-76B on the $5\%$ subset of the ImageNet200 benchmark. Baseline refers to a prompt that includes only the basic elements: a task description and an explanation of the rejection class. As shown in Tab. B.5.4, the inclusion of guidelines and examples results in a nearly $90\%$ reduction in failure cases compared to prompts without them. Examples help address all types of failure cases, while guidelines are especially effective in reducing Failure Case (2), where responses do not follow the required format.
Table B.5.4: Ablation results on the effects of guidelines and examples in reducing failure cases with InternVL2-76B
Prompt (1) (2) (3) (4) (5) (6) Total
Baseline 193 886 - 0 201 0 1,280
+ ID example 18 731 0 0 100 0 849
   + Guidelines 10 493 0 0 128 0 631
     + OoD example 26 34 0 0 79 8 147
@@ B.6 Prompt Format
We provide the full prompt formats used for experiments in Sec. 3.5 and 3.6. In the prompt format, yellow, green, red, blue boxes indicate a task description, an explanation of the reject class, guidelines, and examples for response format, respectively.
@@
Figure B.6.1: Prompt used for experiments in Sec. 3.5
@@ B.6.2 Prompt for Reasoning
Figure B.6.2: Prompt used for reasoning in Sec. 3.6
@@ B.6.3 Prompt for Class order
Figure B.6.3: Prompt used for GPT-4o to group classes for class order in Sec. 3.6
@@ C.1 Suggested Class Post-processing
Despite the guidelines provided in Stage 1, responses that deviated from the guidelines are observed. These failure cases can be broadly categorized into four cases: 1) inclusion of ID class names, 2) failure to suggest exactly $N$ classes for each near- and far-OoD group, 3) duplicated classes, and 4) no suggestion due to failure in recognizing the image. We post-process the suggested answers to refine them into a feasible set of auxiliary OoD classes, except for Case 4. For Case 1, we exclude suggestions that are character-level identical to ID classes. For Case 2, if the total number of suggestions exceeds $2N$, we sample only $2N$ classes from the suggestions. If the total number of suggestions is less than $2N$, we use the suggestions as is without any supplementation. For Case 3, we remove one of duplicated classes.
@@
Valid ACC ($\uparrow$) ECE ($\downarrow$) AURC ($\downarrow$) FPR@90%TPR ($\downarrow$) / FPR@95%TPR ($\downarrow$) / AUROC ($\uparrow$)
InternVL2-26B 61.01 (2,544) 91.23 8.12 92.27 82.73 / 82.73 / 58.31 94.34 / 94.34 / 52.51 38.03 / 38.03 / 80.66 28.86 / 28.86 / 85.25 47.91 / 47.91 / 75.72 73.12 / 73.12 / 63.11
+ GPT-text 69.88 (2,914) 89.58 11.74 82.48 69.44 / 69.44 / 62.17 85.65 / 85.73 / 53.55 26.82 / 28.00 / 84.72 29.20 / 29.20 / 83.51 39.39 / 39.39 / 78.10 62.41 / 62.64 / 65.88
+ ReGuide 86.14 (3,592) 93.53 11.94 55.63 22.39 / 22.89 / 86.53 15.21 / 15.21 / 90.41 1.39 / 1.39 / 98.02 3.93 / 3.93 / 97.05 2.04 / 2.04 / 97.68 10.24 / 10.27 / 93.19
InternVL2-76B 97.26 (4,056) 89.09 5.97 79.87 51.28 / 51.28 / 71.89 71.02 / 71.02 / 62.02 2.20 / 2.20 / 96.43 10.76 / 10.76 / 92.15 14.27 / 14.27 / 90.40 44.46 / 44.46 / 75.30
+ReGuide 95.80 (3,995) 90.93 1.54 9.28 8.05 / 56.36 / 91.35 14.58 / 66.65 / 87.65 0.00 / 59.75 / 95.35 4.08 / 60.00 / 93.38 2.02 / 65.46 / 93.95 8.92 / 64.36 / 90.60
GPT-4o 87.58 (3,652) 90.64 2.25 10.14 8.57 / 14.76 / 93.96 29.25 / 34.50 / 82.28 0.81 / 1.83 / 98.11 5.60 / 6.47 / 95.37 1.21 / 3.63 / 97.82 15.62 / 19.34 / 89.85
+ ReGuide 79.57 (3,318) 91.59 1.54 8.57 0.49 / 18.72 / 96.76 7.53 / 31.17 / 92.56 0.00 / 17.05 / 97.08 1.32 / 26.43 / 95.96 0.15 / 19.66 / 96.82 4.02 / 25.66 / 94.61
@@
As in Sec. B.2, we evaluate the effectiveness of ReGuide on the shared valid query set. As shown in Tab. C.2.2, the overall performance trends among the LVLMs remained consistent with their performance on their own valid query sets. ReGuide enhances OoDD performance across all models, consistent with the results observed on their own valid query sets. We additionally evaluate their FPR at a TPR threshold of $90\%$, considering the high sensitivity to TPR thresholds, as discussed in Sec. 3.6. At the $90\%$ TPR threshold, FPR is significantly reduced, and ReGuide consistently demonstrates lower FPR compared to the baseline across all OoD datasets.
Table C.2.2: ReGuide results on the shared valid query set of the 5% ImageNet200 benchmark. $\downarrow$ and $\uparrow$ mean that lower and higher values are better, respectively. The AURC values are multiplied by $10^{3}$, and the other values are percentages. Bold represents the best performance among the results from each model.
Models ID Near-OoD Far-OoD All OoD
ImageNet200 NINCO SSB-Hard iNaturalist Textures Openimage-O
Valid ACC ($\uparrow$) ECE ($\downarrow$) AURC ($\downarrow$) FPR@90%TPR ($\downarrow$) / FPR@95%TPR ($\downarrow$) / AUROC ($\uparrow$)
InternVL2-26B 93.30 6.15 61.82 86.67 / 86.67 / 56.39 93.78 / 93.78 / 52.83 34.03 / 34.03 / 82.70 31.52 / 31.52 / 83.96 47.20 / 47.20 / 76.12 70.34 / 70.34 / 64.55
+ ReGuide 96.09 9.22 39.58 33.33 / 33.33 / 81.75 17.73 / 17.88 / 90.21 2.09 / 2.09 / 98.75 5.43 / 6.52 / 96.91 2.00 / 2.00 / 98.84 12.23 / 12.39 / 93.22
InternVL2-76B 34.29 92.74 5.03 71.12 65.33 / 65.33 / 66.22 76.98 / 76.98 / 60.39 1.05 / 1.05 / 98.36 10.87 / 10.87 / 93.45 18.40 / 18.40 / 89.68 48.12 / 48.12 / 74.82
+ReGuide (1,430) 93.85 1.40 9.17 6.67 / 56.00 / 93.27 19.75 / 73.09 / 86.39 0.00 / 64.92 / 96.46 6.52 / 69.57 / 93.14 2.80 / 60.40 / 95.10 11.59 / 68.03 / 90.58
GPT-4o 93.85 1.82 2.92 8.00 / 9.33 / 93.86 32.81 / 34.68 / 80.98 1.05 / 1.05 / 97.92 8.70 / 8.70 / 93.66 1.60 / 1.60 / 97.73 18.47 / 19.50 / 88.62
+ ReGuide 92.18 1.39 11.22 1.33 / 28.00 / 96.42 9.80 / 42.77 / 91.41 0.00 / 25.13 / 97.13 3.26 / 33.70 / 95.12 0.40 / 33.20 / 96.62 5.44 / 36.61 / 93.90
Tab. C.2.3 and C.2.4 present example responses of each models in Tab. 4. The first row in Tab. C.2.3 indicates the classes suggested by GPT-4o when ID class names are provided (i.e., GPT-text). The prompt used to obtain class suggestions of GPT-text can be found in Sec. C.5. A class name below each image is the ground truth of the corresponding image. The class suggestion set in GPT-text remains fixed regardless of the input images, as ID classes are static information, whereas ReGuide’s suggestions change based on the input image due to its image-adaptive nature.
The results on InternVL2-26B/-76B and GPT-4o demonstrate the effectiveness of ReGuide. Firstly, ReGuide suggests helpful classes for OoDD. For example, as shown in Tab. C.2.3, InternVL2+ReGuide suggests near-OoD classes for dotted image, such as polka dot, cherry red, or salmon-colored, which are highly close to its ground-truth. These suggested classes successfully guide the model to classify OoD inputs into one of the auxiliary OoD classes. For softball image, the prediction of ReGuide is also closest to the ground-truth.
Table C.2.3: Example responses of InternVL2-26B on $5\%$ subset of ImageNet200 benchmark
@@
Table C.2.4: Example responses of GPT-4o on $5\%$ subset of ImageNet200 benchmark